A car moving with speed v on a straight road can be stopped with in distance d on applying brakes. If same car is moving with speed 3v and brakes provide half retardation,then car will stop after travelling distance?

Grade:11

1 Answers

Manish Kumar
6 years ago
1st speed = vStopped in distance = d2nd speed =3vRetardation =1/2What is distance covered by car while braking or stopping?By puting 3rd law of motion i.e v^2 = u^2+2asSo a=v^2÷2da=1÷2(v^2÷2d)=v^2÷4dd=(3v)^2÷2d=9v^2÷2v^2÷4d=9×4d÷2= 18d